# Quantization Error Propagation: Revisiting Layer-Wise Post-Training Quantization

### v2

@@ A.2 Quantization Error Accumulation
Proposition A.2.
Suppose each activation function $\sigma_{l}$ is $\gamma_{l}$-Lipschitz continuous with respect to the Frobenius norm:
$\forall{\bm{U}},{\bm{V}}\in\mathbb{R}^{n_{l}\times m},\quad\|\sigma_{l}({\bm{U}})-\sigma_{l}({\bm{V}})\|_{F}\leq\gamma_{l}\|{\bm{U}}-{\bm{V}}\|_{F},~{}~{}\gamma_{l}\geq 1.$ (20)
Then, the final activation mismatch ${\bm{\delta}}_{L}={\bm{X}}_{L}-\widehat{{\bm{X}}}_{L}$ satisfies the following bound:
$\left\|{\bm{\delta}}_{L}\right\|_{F}\leq(L-1)\left(\prod_{l=1}^{L-1}\gamma_{l}\|{\bm{W}}_{l}\|_{2}\right)\max_{1\leq k\leq L-1}\frac{\|{\bm{E}}_{k}\|_{2}}{\|{\bm{W}}_{k}\|_{2}}\,\|{\bm{X}}\|_{F}+{\mathcal{O}}\left(\|{\bm{E}}_{1}\|_{F}^{2},\dots,\|{\bm{E}}_{L-1}\|_{F}^{2}\right).$ (21)
@@
We start by expanding the activation difference explicitly:
${\bm{\delta}}_{l+1}={\bm{X}}_{l+1}-\widehat{{\bm{X}}}_{l+1}=\sigma_{l}({\bm{W}}_{l}{\bm{X}}_{l})-\sigma_{l}(\widehat{{\bm{W}}}_{l}\widehat{{\bm{X}}}_{l}).$ (22)
By Lipschitz continuity,
$\|{\bm{\delta}}_{l+1}\|_{F}$ $\leq\gamma_{l}\|{\bm{W}}_{l}{\bm{X}}_{l}-\widehat{{\bm{W}}}_{l}\widehat{{\bm{X}}}_{l}\|_{F}.$ (23)
Explicitly introduce and split terms clearly:
${\bm{W}}_{l}{\bm{X}}_{l}-\widehat{{\bm{W}}}_{l}\widehat{{\bm{X}}}_{l}=-{\bm{E}}_{l}{\bm{X}}_{l}+\widehat{{\bm{W}}}_{l}{\bm{\delta}}_{l}.$ (24)
Hence,
$\|{\bm{\delta}}_{l+1}\|_{F}\leq\gamma_{l}(\|{\bm{E}}_{l}{\bm{X}}_{l}\|_{F}+\|\widehat{{\bm{W}}}_{l}{\bm{\delta}}_{l}\|_{F}).$ (25)
Using sub-multiplicativity of the Frobenius norm and discarding the higher-order term, one obtains the recursive inequality
$\|{\bm{\delta}}_{l+1}\|_{F}\leq\gamma_{l}\left(\|{\bm{E}}_{l}\|_{2}\|{\bm{X}}_{l}\|_{F}+\|\widehat{{\bm{W}}}_{l}\|_{2}\|{\bm{\delta}}_{l}\|_{F}\right).$ (26)
Recursively,
$\|{\bm{X}}_{2}\|_{F}$ $=\|\sigma_{1}({\bm{W}}_{1}{\bm{X}}_{1})\|_{F}\leq\gamma_{1}\|{\bm{W}}_{1}\|_{2}\|{\bm{X}}\|_{F},$ (27)
$\|{\bm{X}}_{3}\|_{F}$ $\leq\gamma_{2}\|{\bm{W}}_{2}\|_{2}\|{\bm{X}}_{2}\|_{F}\leq\gamma_{2}\gamma_{1}\|{\bm{W}}_{2}\|_{2}\|{\bm{W}}_{1}\|_{2}\|{\bm{X}}\|_{F},$ (28)
and inductively
$\|{\bm{X}}_{l}\|_{F}\leq\left(\prod_{s=1}^{l-1}\gamma_{s}\|{\bm{W}}_{s}\|_{2}\right)\|{\bm{X}}\|_{F},~{}~{}l\geq 1.$ (29)
Insert (26),
$\|{\bm{\delta}}_{2}\|_{F}$ $\leq\gamma_{1}\|{\bm{E}}_{1}\|_{2}\|{\bm{X}}\|_{F},$ (30)
$\|{\bm{\delta}}_{3}\|_{F}$ $\leq\gamma_{2}\|{\bm{E}}_{2}\|_{2}\gamma_{1}\|{\bm{W}}_{1}\|_{2}\|{\bm{X}}\|_{F}+\gamma_{2}\|{\bm{W}}_{2}\|_{2}\gamma_{1}\|{\bm{E}}_{1}\|_{2}\|{\bm{X}}\|_{F}.$ (31)
The pattern is
$\|{\bm{\delta}}_{l+1}\|_{F}\leq\sum_{k=1}^{l}\left(\prod_{s=k+1}^{l}\gamma_{s}\|{\bm{W}}_{s}\|_{2}\right)\gamma_{k}\|{\bm{E}}_{k}\|_{2}\left(\prod_{t=1}^{k-1}\gamma_{t}\|{\bm{W}}_{t}\|_{2}\right)\|{\bm{X}}\|_{F}.$ (32)
Taking $l=L-1$,
$\|{\bm{\delta}}_{L}\|_{F}\leq\sum_{k=1}^{L-1}\frac{\displaystyle\prod_{l=1}^{L-1}\gamma_{l}\|{\bm{W}}_{l}\|_{2}}{\|{\bm{W}}_{k}\|_{2}}\;\gamma_{k}\|{\bm{E}}_{k}\|_{2}\,\|{\bm{X}}\|_{F}+{\mathcal{O}}\left(\|{\bm{E}}_{1}\|_{F}^{2},\dots,\|{\bm{E}}_{L-1}\|_{F}^{2}\right).$ (33)
Define
$a_{k}\coloneqq\dfrac{\gamma_{k}\|{\bm{E}}_{k}\|_{2}}{\|{\bm{W}}_{k}\|_{2}}.$ (34)
Since $\sum_{k=1}^{L-1}a_{k}\leq(L-1)\max_{k}a_{k}$, inequality (33) implies
$\|{\bm{\delta}}_{L}\|_{F}\leq(L-1)\,\left(\prod_{l=1}^{L-1}\gamma_{l}\|{\bm{W}}_{l}\|_{2}\right)\max_{1\leq k\leq L-1}\frac{\|{\bm{E}}_{k}\|_{2}}{\|{\bm{W}}_{k}\|_{2}}\|{\bm{X}}\|_{F}+{\mathcal{O}}\left(\|{\bm{E}}_{1}\|_{F}^{2},\dots,\|{\bm{E}}_{L-1}\|_{F}^{2}\right).$ (35)
Proposition A.3.
For every hidden layer $l\in\{1,\dots,L-1\}$ let
$\gamma_{l}\|{\bm{W}}_{l}\|_{2}\geq 1+\varepsilon,~{}~{}\varepsilon>0.$ (36)
Suppose that the calibration input and the weight errors are uniformly bounded and non–degenerate,
$\|{\bm{X}}\|_{F}\leq C,~{}~{}c_{E}\leq\|{\bm{E}}_{k}\|_{2}\leq M,~{}~{}k=1,\dots,L-1,$ (37)
where $C,M,c_{E}$ are constants independent of the depth $L$. Then
${\bm{G}}_{L}\coloneqq\prod_{l=1}^{L}\gamma_{l}\|{\bm{W}}_{l}\|_{2}\geq(1+\varepsilon)^{L},$ (38)
and the first–order component of the activation mismatch ${\bm{\delta}}_{L}\coloneqq{\bm{X}}_{L}-\widehat{{\bm{X}}}_{L}$ satisfies
$\|{\bm{\delta}}_{L}\|_{F}\geq c_{\ast}(1+\varepsilon)^{L-1},~{}~{}c_{\ast}=\frac{\varepsilon}{2+\varepsilon}c_{E}C.$ (39)
Consequently
$\|{\bm{\delta}}_{L}\|_{F}=\Omega\left((1+\varepsilon)^{L}\right)$ (40)
as $L\to\infty$.
@@
By (37), $\|{\bm{E}}_{l}{\bm{X}}_{l}\|_{F}\geq c_{E}\|{\bm{X}}_{l}\|_{F}$ and $\|{\bm{X}}_{l}\|_{F}\geq\|{\bm{X}}\|_{F}\geq C$, thus
$\|{\bm{E}}_{l}{\bm{X}}_{l}\|_{F}\geq c_{\!E}C.$ (41)
Because $\gamma_{l}\|{\bm{W}}_{l}\|_{2}\geq 1+\varepsilon$,
$\|{\bm{W}}_{l}\|_{2}\|{\bm{\delta}}_{l}\|_{F}\leq\frac{\|{\bm{\delta}}_{l}\|_{F}}{1+\varepsilon}.$ (42)
Combine these:
@@
$\|{\bm{\delta}}_{l+1}\|_{F}\geq\gamma_{l}\left(c_{E}C-\frac{\|{\bm{\delta}}_{l}\|_{F}}{1+\varepsilon}\right).$ (43)
We use induction. When $l=1$, the following inequality holds:
$\|{\bm{\delta}}_{2}\|_{F}\geq\gamma_{1}\left(c_{E}C-\frac{c_{*}}{1+\varepsilon}\right)=\frac{\varepsilon}{2+\varepsilon}\gamma_{1}c_{E}C\geq(1+\varepsilon)c_{\ast}.$ (44)
Assume $\|{\bm{\delta}}_{l}\|_{F}\geq(1+\varepsilon)^{l-1}c_{\ast}$. Then
$\frac{\|{\bm{\delta}}_{l}\|_{F}}{1+\varepsilon}\leq(1+\varepsilon)^{l-2}c_{\ast}\leq\frac{2+\varepsilon}{\varepsilon}c_{\ast},$ (45)
so the bracket in (43) is at least $c_{\ast}$ and
$\|{\bm{\delta}}_{l+1}\|_{F}\geq\gamma_{l}c_{\ast}\geq(1+\varepsilon)c_{\ast},$ (46)
which equals $(1+\varepsilon)^{l}c_{\ast}$. By induction (39) holds. Because $c_{\ast}>0$ is independent of $L$, (39) is equivalent to the asymptotic lower bound
$\|{\bm{\delta}}_{L}\|_{F}=\Omega\left((1+\varepsilon)^{L}\right).$ (47)
A.3 Derivation of Theorem 5.2 and Corollary 5.4
This section provides a rigorous statement and proof of Theorem 5.2 and Corollary 5.4. We first restate Theorem 5.2 formally bellow.
Theorem A.4.
Consider an $L$-layer neural network defined by
$f_{{\bm{\theta}}}({\bm{X}})=\sigma_{L}({\bm{W}}_{L}\sigma_{L-1}({\bm{W}}_{L-1}\cdots\sigma_{2}({\bm{W}}_{2}\sigma_{1}({\bm{W}}_{1}{\bm{X}}))\cdots)).$ (48)
where ${\bm{\theta}}$ denotes the set of all full-precision parameters $\{{\bm{W}}_{l}\}_{l=1}^{L}$. Assume each activation function $\sigma_{l}$ is $\gamma_{l}$-Lipschitz with respect to the Frobenius norm:
[…]
For layer-wise PTQ, let the quantized weights obtained via the standard method (BASE) and the proposed QEP method as
$\widehat{{\bm{W}}}_{l}^{\mathrm{BASE}}$ $={\bm{W}}_{l}+{\bm{E}}_{l},$ (49)
$\widehat{{\bm{W}}}_{l}^{\mathrm{QEP}}$ $={\bm{W}}_{l}^{\ast}(\alpha_{l})+{\bm{E}}_{l},$ (50)
where ${\bm{E}}_{l}$ represents the quantization error matrix at layer $l$. Under these conditions, the output quantization error from the QEP method is upper-bounded by that of the BASE method:
$\left\|f_{{\bm{\theta}}}({\bm{X}})-f_{\widehat{{\bm{\theta}}}^{\mathrm{QEP}}}({\bm{X}})\right\|_{F}\leq\left\|f_{{\bm{\theta}}}({\bm{X}})-f_{\widehat{{\bm{\theta}}}^{\mathrm{BASE}}}({\bm{X}})\right\|_{F},$ (51)
where $\widehat{{\bm{\theta}}}^{\mathrm{QEP}}$ and $\widehat{{\bm{\theta}}}^{\mathrm{BASE}}$ denote the sets of parameters quantized by QEP and the base PTQ method, respectively.

### v3

@@ B.2 Quantization Error Accumulation
Proposition B.2.
Assume each activation $\sigma_{l}:{\mathbb{R}}^{n_{l}\times m}\to{\mathbb{R}}^{n_{l}\times m}$ is $\gamma_{l}$-Lipschitz with respect to the Frobenius norm and satisfies $\sigma_{l}({\bm{0}})={\bm{0}}$:
$\|\sigma_{l}({\bm{U}})-\sigma_{l}({\bm{V}})\|_{F}\leq\gamma_{l}\|{\bm{U}}-{\bm{V}}\|_{F},~~\gamma_{l}>0.$ (30)
Let ${\bm{X}}_{1}=\widehat{{\bm{X}}}_{1}={\bm{X}}$; for $l=1,\dots,L-1$ define
${\bm{X}}_{l+1}=\sigma_{l}({\bm{W}}_{l}{\bm{X}}_{l}),~~\widehat{{\bm{X}}}_{l+1}=\sigma_{l}(\widehat{{\bm{W}}}_{l}\widehat{{\bm{X}}}_{l}),~~\widehat{{\bm{W}}}_{l}={\bm{W}}_{l}+{\bm{E}}_{l},$ (31)
and ${\bm{\delta}}_{l}\coloneq{\bm{X}}_{l}-\widehat{{\bm{X}}}_{l}$. Assume $\|{\bm{W}}_{l}\|_{2}>0$ for all $l=1,\dots,L-1$ and set
$G_{L-1}\coloneq\prod_{l=1}^{L-1}\gamma_{l}\|{\bm{W}}_{l}\|_{2},~~r\coloneq\max_{1\leq k\leq L-1}\frac{\|{\bm{E}}_{k}\|_{2}}{\|{\bm{W}}_{k}\|_{2}}.$ (32)
Then the final activation mismatch satisfies the explicit bound
$\|{\bm{\delta}}_{L}\|_{F}\leq\bigl((1+r)^{L-1}-1\bigr)G_{L-1}\|{\bm{X}}\|_{F}.$ (33)
@@
Since $\sigma_{l}({\bm{0}})={\bm{0}}$ and $\sigma_{l}$ are $\gamma_{l}$-Lipschitz,
$\|{\bm{X}}_{l+1}\|_{F}=\|\sigma_{l}({\bm{W}}_{l}{\bm{X}}_{l})-\sigma_{l}({\bm{0}})\|_{F}\leq\gamma_{l}\|{\bm{W}}_{l}{\bm{X}}_{l}\|_{F}\leq\gamma_{l}\|{\bm{W}}_{l}\|_{2}\|{\bm{X}}_{l}\|_{F}.$ (34)
By induction,
$\|{\bm{X}}_{l}\|_{F}\leq\left(\prod_{t=1}^{l-1}\gamma_{t}\|{\bm{W}}_{t}\|_{2}\right)\|{\bm{X}}\|_{F}=G_{l-1}\|{\bm{X}}\|_{F},~~l\geq 1.$ (35)
Using Lipschitz continuity again,
$\|{\bm{\delta}}_{l+1}\|_{F}$ $=\|\sigma_{l}({\bm{W}}_{l}{\bm{X}}_{l})-\sigma_{l}(\widehat{{\bm{W}}}_{l}\widehat{{\bm{X}}}_{l})\|_{F}$ (36)
$\leq\gamma_{l}\|{\bm{W}}_{l}{\bm{X}}_{l}-\widehat{{\bm{W}}}_{l}\widehat{{\bm{X}}}_{l}\|_{F}.$ (37)
Since $\widehat{{\bm{W}}}_{l}={\bm{W}}_{l}+{\bm{E}}_{l}$ and ${\bm{\delta}}_{l}={\bm{X}}_{l}-\widehat{{\bm{X}}}_{l}$,
${\bm{W}}_{l}{\bm{X}}_{l}-\widehat{{\bm{W}}}_{l}\widehat{{\bm{X}}}_{l}={\bm{W}}_{l}{\bm{X}}_{l}-({\bm{W}}_{l}+{\bm{E}}_{l})\widehat{{\bm{X}}}_{l}=-{\bm{E}}_{l}{\bm{X}}_{l}+\widehat{{\bm{W}}}_{l}{\bm{\delta}}_{l},$ (38)
hence, by the triangle inequality and $\|AB\|_{F}\leq\|A\|_{2}\|B\|_{F}$,
$\|{\bm{\delta}}_{l+1}\|_{F}\leq\gamma_{l}\Bigl(\|{\bm{E}}_{l}\|_{2}\|{\bm{X}}_{l}\|_{F}+\|\widehat{{\bm{W}}}_{l}\|_{2}\|{\bm{\delta}}_{l}\|_{F}\Bigr).$ (39)
By definition of $r$, $\|{\bm{E}}_{l}\|_{2}\leq r\|{\bm{W}}_{l}\|_{2}$. Additionally, $\|\widehat{{\bm{W}}}_{l}\|_{2}\leq\|{\bm{W}}_{l}\|_{2}+\|{\bm{E}}_{l}\|_{2}\leq(1+r)\|{\bm{W}}_{l}\|_{2}$. Combining these with Eq. (35) in Eq. (39) yields
$\|{\bm{\delta}}_{l+1}\|_{F}\leq\gamma_{l}\|{\bm{W}}_{l}\|_{2}\Bigl(rG_{l-1}\|{\bm{X}}\|_{F}+(1+r)\|{\bm{\delta}}_{l}\|_{F}\Bigr).$ (40)
Define the normalized quantity
$a_{l}\coloneq\frac{\|{\bm{\delta}}_{l}\|_{F}}{G_{l-1}\|{\bm{X}}\|_{F}},~~l\geq 1.$ (41)
Note $a_{1}=0$ because ${\bm{\delta}}_{1}={\bm{0}}$. Dividing the previous inequality by $G_{l}\|{\bm{X}}\|_{F}$ (where $G_{l}=G_{l-1}\gamma_{l}\|{\bm{W}}_{l}\|_{2}$) yields
$a_{l+1}\leq r+(1+r)a_{l}.$ (42)
Let $b_{l}\coloneq a_{l}+1$. Then $b_{l+1}\leq(1+r)b_{l}$ and $b_{1}=1$; hence $b_{l}\leq(1+r)^{l-1}$; therefore
$a_{l}\leq(1+r)^{l-1}-1.$ (43)
Taking $l=L$ gives Eq. (33). ∎
Proposition B.3.
Assume each activation $\sigma_{l}:{\mathbb{R}}^{n_{l}\times m}\to{\mathbb{R}}^{n_{l}\times m}$ is $\gamma_{l}$-Lipschitz with respect to the Frobenius norm and satisfies $\sigma_{l}({\bm{0}})={\bm{0}}$:
$\|\sigma_{l}({\bm{U}})-\sigma_{l}({\bm{V}})\|_{F}\leq\gamma_{l}\|{\bm{U}}-{\bm{V}}\|_{F}.$ (44)
Assume moreover that $\sigma_{l}$ is Fréchet differentiable at the full-precision pre-activation ${\bm{Y}}_{l}\coloneq{\bm{W}}_{l}{\bm{X}}_{l}$ for each $l=1,\dots,L-1$. Let ${\bm{X}}_{1}=\widehat{{\bm{X}}}_{1}={\bm{X}}$ and for $l=1,\dots,L-1$ define
${\bm{X}}_{l+1}=\sigma_{l}({\bm{W}}_{l}{\bm{X}}_{l}),~~\widehat{{\bm{X}}}_{l+1}=\sigma_{l}(\widehat{{\bm{W}}}_{l}\widehat{{\bm{X}}}_{l}),~~\widehat{{\bm{W}}}_{l}={\bm{W}}_{l}+{\bm{E}}_{l}.$ (45)
Assume $\|{\bm{W}}_{k}\|_{2}>0$ for $k=1,\dots,L-1$ and define
${\bm{G}}_{L-1}\coloneq\prod_{l=1}^{L-1}\gamma_{l}\|{\bm{W}}_{l}\|_{2},~~r\coloneq\max_{1\leq k\leq L-1}\frac{\|{\bm{E}}_{k}\|_{2}}{\|{\bm{W}}_{k}\|_{2}}.$ (46)
Define the first-order component ${\bm{\delta}}_{L}^{(1)}$ as follows. For $t\in{\mathbb{R}}$ let $\widehat{{\bm{W}}}_{l}(t)\coloneq{\bm{W}}_{l}+t{\bm{E}}_{l}$ and define $\widehat{{\bm{X}}}_{1}(t)\coloneq{\bm{X}}$, $\widehat{{\bm{X}}}_{l+1}(t)\coloneq\sigma_{l}(\widehat{{\bm{W}}}_{l}(t)\widehat{{\bm{X}}}_{l}(t))$. Let ${\bm{\delta}}_{l}(t)\coloneq{\bm{X}}_{l}-\widehat{{\bm{X}}}_{l}(t)$. Then ${\bm{\delta}}_{L}^{(1)}$ is defined by the derivative
${\bm{\delta}}_{L}^{(1)}\coloneq\left.\frac{d}{dt}\right|_{t=0}{\bm{\delta}}_{L}(t).$ (47)
Under these assumptions,
$\|{\bm{\delta}}_{L}^{(1)}\|_{F}\leq(L-1)r{\bm{G}}_{L-1}\|{\bm{X}}\|_{F}.$ (48)
In particular, if $\gamma_{l}\|{\bm{W}}_{l}\|_{2}\leq 1+\varepsilon$ for all $l$, then
$\|{\bm{\delta}}_{L}^{(1)}\|_{F}\leq(L-1)r(1+\varepsilon)^{L-1}\|{\bm{X}}\|_{F}.$ (49)
@@
Since $\sigma_{l}({\bm{0}})={\bm{0}}$ and $\sigma_{l}$ is $\gamma_{l}$-Lipschitz,
$\|{\bm{X}}_{l+1}\|_{F}=\|\sigma_{l}({\bm{W}}_{l}{\bm{X}}_{l})-\sigma_{l}({\bm{0}})\|_{F}\leq\gamma_{l}\|{\bm{W}}_{l}{\bm{X}}_{l}\|_{F}\leq\gamma_{l}\|{\bm{W}}_{l}\|_{2}\|{\bm{X}}_{l}\|_{F}.$ (50)
By induction,
$\|{\bm{X}}_{l}\|_{F}\leq\left(\prod_{t=1}^{l-1}\gamma_{t}\|{\bm{W}}_{t}\|_{2}\right)\|{\bm{X}}\|_{F},~~l\geq 1.$ (51)
Fix $l\in\{1,\dots,L-1\}$. By assumption, $\sigma_{l}$ is Fréchet differentiable at ${\bm{Y}}_{l}\coloneq{\bm{W}}_{l}{\bm{X}}_{l}$. Let ${\bm{J}}_{l}\coloneq D\sigma_{l}({\bm{Y}}_{l})$ denote its Fréchet derivative. Because $\sigma_{l}$ is $\gamma_{l}$-Lipschitz, the operator norm of ${\bm{J}}_{l}$, induced by the Frobenius norm, satisfies
$\|{\bm{J}}_{l}\|_{\mathrm{op}}\leq\gamma_{l}.$ (52)
Indeed, for any $\bm{H}$},
$\|{\bm{J}}_{l}[{\bm{H}}]\|_{F}=\lim_{t\to 0}\frac{\|\sigma_{l}({\bm{Y}}_{l}+t{\bm{H}})-\sigma_{l}({\bm{Y}}_{l})\|_{F}}{|t|}\leq\lim_{t\to 0}\frac{\gamma_{l}\|t{\bm{H}}\|_{F}}{|t|}=\gamma_{l}\|{\bm{H}}\|_{F}.$ (53)
Now consider $\widehat{{\bm{X}}}_{l+1}(t)=\sigma_{l}(\widehat{{\bm{W}}}_{l}(t)\widehat{{\bm{X}}}_{l}(t))$. Since matrix multiplication is smooth and $\sigma_{l}$ is differentiable at ${\bm{Y}}_{l}$, the chain rule gives
$\left.\frac{d}{dt}\right|_{t=0}\widehat{{\bm{X}}}_{l+1}(t)={\bm{J}}_{l}\!\left[{\bm{E}}_{l}{\bm{X}}_{l}+{\bm{W}}_{l}\left.\frac{d}{dt}\right|_{t=0}\widehat{{\bm{X}}}_{l}(t)\right].$ (54)
@@
Because ${\bm{\delta}}_{l}(t)={\bm{X}}_{l}-\widehat{{\bm{X}}}_{l}(t)$, we have $\left.\frac{d}{dt}\right|_{t=0}\widehat{{\bm{X}}}_{l}(t)=-{\bm{\delta}}_{l}^{(1)}$. Therefore,
${\bm{\delta}}_{l+1}^{(1)}=-{\bm{J}}_{l}[{\bm{E}}_{l}{\bm{X}}_{l}-{\bm{W}}_{l}{\bm{\delta}}_{l}^{(1)}]=-{\bm{J}}_{l}({\bm{E}}_{l}{\bm{X}}_{l})+{\bm{J}}_{l}({\bm{W}}_{l}{\bm{\delta}}_{l}^{(1)}),~~{\bm{\delta}}_{1}^{(1)}={\bm{0}}.$ (55)
Taking Frobenius norms, using Eq. (52) and $\|{\bm{A}}{\bm{B}}\|_{F}\leq\|{\bm{A}}\|_{2}\|{\bm{B}}\|_{F}$,
$\|{\bm{\delta}}_{l+1}^{(1)}\|_{F}\leq\gamma_{l}\|{\bm{E}}_{l}\|_{2}\|{\bm{X}}_{l}\|_{F}+\gamma_{l}\|{\bm{W}}_{l}\|_{2}\|{\bm{\delta}}_{l}^{(1)}\|_{F}.$ (56)
Define $a_{l}\coloneq\|{\bm{\delta}}_{l}^{(1)}\|_{F}$. From Eq. (56) and $a_{1}=0$, straightforward induction yields
$a_{L}\leq\sum_{k=1}^{L-1}\left(\prod_{s=k+1}^{L-1}\gamma_{s}\|{\bm{W}}_{s}\|_{2}\right)\gamma_{k}\|{\bm{E}}_{k}\|_{2}\|{\bm{X}}_{k}\|_{F}.$ (57)
Apply Eq. (51) to $\|{\bm{X}}_{k}\|_{F}$ and factor out ${\bm{G}}_{L-1}$:
$\|{\bm{\delta}}_{L}^{(1)}\|_{F}\leq{\bm{G}}_{L-1}\left(\sum_{k=1}^{L-1}\frac{\|{\bm{E}}_{k}\|_{2}}{\|{\bm{W}}_{k}\|_{2}}\right)\|{\bm{X}}\|_{F}\leq(L-1)r{\bm{G}}_{L-1}\|{\bm{X}}\|_{F},$ (58)
which is Eq. (48). If additionally $\gamma_{l}\|{\bm{W}}_{l}\|_{2}\leq 1+\varepsilon$ for all $l$, then ${\bm{G}}_{L-1}\leq(1+\varepsilon)^{L-1}$, proving Eq. (49). ∎
Proposition B.4.
Consider the 1-dimensional network ($d_{l}=n_{l}=1$) with $\sigma_{l}(z)=z$. Let $W_{l}=1+\varepsilon$ for all $l=1,\dots,L-1$ with $\varepsilon>0$, and let the input be $X=C>0$. Choose quantized weights
$\widehat{W}_{l}=W_{l}+E_{l},~~E_{l}\equiv c_{E}>0,~~l=1,\dots,L-1.$ (59)
Then, for all $L\geq 2$, the exact activation mismatch at layer $L$ satisfies
$|\delta_{L}|=|X_{L}-\widehat{X}_{L}|\geq(L-1)c_{E}C(1+\varepsilon)^{L-2}.$ (60)
In particular,
$|\delta_{L}|\geq\frac{c_{E}C}{1+\varepsilon}(1+\varepsilon)^{L-1}.$ (61)